| Operation Summary for December 5, 2011 - December 7, 2011 |
|
Monday December 5 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with NuMI, MiniBooNE, and Meson FTBFT-978 experiments taking beam. During the day shift, operators turned on the Muon Rings, Transport, Pretarget, and Prevault at 9:30 AM in preparation for studies. At 9:33 AM, the beam permit tripped off for the P1 and P2 transfer lines when a Lambertson (LAM52) tripped off on a load transfer switch fault; everything reset. Mechanical technicians began transferring gas out of the Pelletron at 9:47 AM. A Muon Rings person began a 2.5-hour study at 9:56 AM. A MI person began a one-hour study at 10:52 AM. Operators rebooted a NuMI BPM VME (NBPM6N) at 11:33 AM. Operators searched and secured the D-Zero collision hall at 2:16 PM. Operators prepared the Muon Rings, Transport, Pretarget, and Prevault area for access at 2:30 PM. During the evening shift, Operations disabled T-978 beam at 7 PM. At 7:40 PM, the NuMI beam permit tripped off when a NuMI BPM (NBPM65) stopped polling; everything reset. Tuesday December 6 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, operators completed a Booster injection tune at 1:18 AM. Operators loaded protons into the Recycler at 1:33 AM for studies. Operators reported at 3:10 AM that the Linac Klystron distribution water skid LCW temperature had gone up; operators manually adjusted a valve to lower the temperature. Operations established beam to T-978 at 7:03 AM. During the day shift, Muon Ring personnel began studies. Operators reported at 1:26 PM that a Booster kicker (B:MSK11) wasn’t working properly, which was holding off beam to T-978 and causing beam position problems for Muon Rings studiers. Booster kicker personnel had the kicker repaired and back on line by 2:14 PM. At 2:50 PM, a MI RF station (MIRF1) tripped off and wouldn’t reset; RF personnel investigated and reported that the modulator had failed. MI will run without this station. During the evening shift, the Muon Ring studies ended at 5:26 PM and operators prepared the area for access. Operations disabled the T-978 at 7:33 PM; this ended their experimental run. A new MTest experiment will begin on Thursday, 12/8/11. Operators completed a Booster injection tune at 7:39 PM. Operators rebooted a NuMI BPM (NBPM6N) at 9:15 PM. At 10:33 PM, two I- Source ion pump power supplies (#1 and 3) tripped off; operators reset them. At 11:54 PM, operators reported that the I- Source arc current began to fall; operators increased the arc supply voltage. Wednesday December 7 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, operators completed a MI tune at 12:06 AM. The NuMI beam permit tripped off at 1:19 AM due to an incomplete gap fault, but no MI RF stations had tripped off; it reset. At 5:43 AM, a Booster RF station (BRF11) tripped off for the first of three times in about 30 minutes on a ground fault indication; it reset. Operators prepared the Transfer Hall for access at 7:46 AM. Operators reported short downtimes due to problems with LRF2 and LRF4. Operations Coordinator (OpsCo) Plans for Today The OpsCo approved of the plans by Booster, MI, and Muon personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 12/07/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |